OpenStudy (anonymous):

limx->0 10sin(5x)/2x=25

hartnn (hartnn):

do you know what \(\Large \lim \limits_{x \rightarrow 0}\dfrac{\sin x}{x}=...?\)

hartnn (hartnn):

 did you get this adjustment ?